# Textual Entailment and Token Probability as Bias Evaluation Metrics

### v1

@@ 3.1 WQ-NLI Baseline Results
Table 3: Comparison of TP and NLI bias scores for raw and debiased models. NLI bias scores are measured by $M_{1}$ in this table. TP Raw, TP News, and TP Twitter columns are reproduced from results in Felkner et al. (2023).
3.2 Correlation Between NLI and Token Probability Bias Scores
In this section, we consider the overall bias scores for six models (four BERT variants and two RoBERTa variants), each tested under three debiasing conditions: off-the-shelf, continued pretraining on news, and on continued pretraining on Twitter data from the affected community. In general, we see very poor correlation between percentile bias scores calculated from token probability and those calculated from various NLI metrics. The maximum $R^{2}$ value found is $0.328$. From this analysis we conclude that token probability and NLI, when formulated as percentile bias scores, do not seem to be measuring the same model behavior.
3.3 Correlation Among NLI Bias Metrics
When comparing token probability and NLI bias scores, we notice that the overall behavior of NLI bias metrics is concerningly random. None of the metrics behave predictably with respect to token probability bias scores. We want to understand if this is caused by the relationship between TP and NLI bias scores, or if the NLI metrics we tested are generally brittle with high randomness. Out of 28 pairwise combinations of NLI metrics, we find only five with $R^{2}>0.5$. These pairs are listed in Table 4.
Metric Pair $R^{2}$
$M_{1}$ vs. $M_{6}$ $0.567$
$M_{2}$ vs. $M_{3}$ $0.633$
$M_{3}$ vs. $M_{8}$ $0.852$
$M_{2}$ vs. $M_{8}$ $0.558$
$M_{5}$ vs. $M_{6}$ $0.955$
Table 4: NLI metric pairs with high $R^{2}$ values
Of these pairs, $M_{2}$ and $M_{3}$, as well as $M_{5}$ and $M_{6}$, are very similarly formulated, so those correlations are unsurprising. The most interesting result here is the relatively strong correlation between $M_{3}$ and $M_{8}$, which are defined very differently but behave similarly. This is weak evidence that either probability thresholds or argmax defintions might be the most useful conversions from NLI classification probabilities to percentile bias scores. However, the overall behavior of the tested NLI metrics seems to be brittle and hard to predict. Given this finding, we next attempt to determine whether there is a meaningful difference between raw TP and NLI scores, or whether the issue is at the level of conversion to percentile bias scores.
3.4 Per-Category Analysis of TP and NLI Results
Until this point, we have been considering percentile bias scores that are calculated based on comparing the raw scores for the stereotypical and counterstereotypical sides of the WQ-TP/WQ-NLI datasets. In the next few sections, we will consider the raw scores themselves. For token probability bias evaluation, the raw score for each test sentence is the sum of pseudo-log-probabilities for each of the unmodified tokens ($U$), i.e. tokens that are shared between the stereotypical and counterstereotypical test sentences, while holding the modified tokens ($M$) constant:
$\text{score}(S)=\sum_{u\in U}\log P(u|U_{\backslash u},M,\theta)$ (1)
[…]
Figure 1: Entailment and token probability bias scores for selected categories on three BERT Large Uncased models (undebiased, news-debiased, and twitter-debiased). These categories appear to be adequately debiased on the token probability metric, but the NLI metric shows lingering model bias.
We find that, in general, token probability bias metrics are more sensitive to debiasing efforts, while NLI bias metrics seem more likely to detect areas where models are insufficiently debiased. The two metrics tend to be closer in scores for larger models, whereas differences are more distinct for smaller model variants. Across debiased models, we find that two predicate categories are consistently “underdebiased”: drug use and naturalness/normalness. We also observe similar, but less pronounced, behavior for the sensitivity/emotion/attention-seeking category. NLI score differences for these predicate categories are consistently significantly above median NLI score differences in almost all cases, while TP score differences for these categories are at or below median TP score differences. Thus, we conclude that NLI bias evaluation is more able to identify which stereotypes are insufficiently covered in model debiasing than TP evaluation. Figure 1 shows these results for BERT Large Uncased; results for other models can be found in Appendix A.
We observe more mixed patterns when breaking down test instances by bias target group. There is considerable variance by model and finetuning condition. Token probability is still more sensitive to debiasing than NLI; however, there is no clear pattern of which evaluation task was more likely to detect “underdebiasing” for specific demographic groups. Token probability seems to be slightly more likely to detect underdebiasing on the worse-performing news debiasing data, while NLI seems more likely to detect underdebiasing on the better-performing Twitter debiasing dataset. This is an indication, albeit a weak one, that NLI may be more sensitive to small, specific issues in models that otherwise appear to be relatively unbiased.
@@ 3.5 Mutual Information Analysis
We want to understand why the the log-prob differences in the previous analysis are so spread out, as represented by the long whiskers on the boxplots in Figure 1 and related figures in the Appendices. We first conduct an exploratory analysis of extreme cases for all models. For each model and finetuning condition combination (27 total), we identify the 500 largest and smallest log-prob differences for both TP and NLI. We then manually examine these extreme examples for trends.
@@
We notice that a very small number of specific predicates dominate the most extreme examples; however, these predicates vary by model without a clear pattern across models. We also notice that several models seem to be very sensitive to the choice of counterfactual identity, with “cis” particularly overrepresented in highest-difference cases. This is likely because the tested models have seen relatively little training data using the word “cis,” and more data containing other counterfactual identities “straight,” “heterosexual,” and “cisgender.” In the analysis of extreme examples, we observe that the changes between raw and debiased models are largely as we expected. Debiasing reduces the magnitude of extreme differences and makes the extremes less dominated by specific predicates and counterfactuals.
@@
From these results, it is clear that there is a lot of noise in bias scores (both percentile bias scores and raw log-probabilities) at both the per-instance and per-model levels. We thus conduct a more systematic analysis of which specific word choices have the largest effects on log-probability differences. We treat this as a feature selection problem, in which the potential features are binary columns corresponding to each possible model, finetuning condition, predicate, predicate category, name, pronoun, bias target group, counterfactual identity, and sentence template. We use mutual information regression, where these binary features predict a difference in log-probabilities, in order to determine the most sensitive, and thus meaningful, features.
@@
In this analysis, we expect to see higher mutual information for factors that should affect bias scores: model, finetuning condition, bias target groups, and predicate categories. We also expect that specific predicates will have some impact, but extremely high MI for certain predicates indicates that they may be introducing noise into the evaluation. For counterfactuals, names, pronouns, and templates, we expect to see very low MI values. Higher MI values for these categories indicate that incidental wording choices are introducing significant noise into the bias evaluation. The top 15 MI factors for both TP and NLI evaluation are listed in Table 5.
TP Factor TP MI Value NLI Factor NLI MI Value
raw 0.033 gpt2_xl 0.037
Straight 0.025 pred cat: naturalness/normalness 0.037
twitter 0.025 predicate: are deviant 0.026
Cisgender 0.015 Straight 0.023
Heterosexual 0.015 LGBTQ 0.017
gpt2_medium 0.014 predicate: are sexually deviant 0.016
subject_is_and 0.009 Cis 0.015
Cis 0.008 Transgender 0.015
Bisexual 0.007 roberta_base 0.012
gpt2 0.007 bert_base_cased 0.010
no_predicate 0.007 bert_base_uncased 0.010
Asexual 0.006 pred cat: sexual practices 0.010
Queer 0.006 Queer 0.010
LGBTQ 0.006 pred cat: lack of belonging 0.010
gpt2_xl 0.005 Cisgender 0.009
@@
Table 5: Factors with highest mutual information values for token probability (left) and NLI (right). Specific wording of counterfactuals has a larger impact that expected.
@@
As in the previous section, we find that token probability evaluation is more sensitive to debiasing than NLI. In particular, the token probability bias evaluation seems to be much more sensitive to choice of counterfactuals than NLI evaluation. This means that the wording of counterstereotypical sentences could unintentionally skew the resulting TP bias scores. NLI seems to be more sensitive to specific predicates and predicate categories. This means that NLI is likely better at detecting stereotypes where the model’s latent associations are strongest. However, the increased sensitivity to specific predicates indicates that wording choices of bias definitions are more likely to introduce noise into NLI bias evaluation.
@@ 4.1 Bias Measurement in LLMs
Defining what constitutes “social bias” in language model behavior is non-trivial, but a clear definition of “bias” is necessary for meaningful discussion on bias measurement and mitigation Blodgett et al. (2020). In this work, we use the definition from Gallegos et al. (2024): “disparate treatment or outcomes between social groups that arise from historical and structural power asymmetries,” which includes both representational and allocational harms. In order to detect social bias, language models are tested on benchmark datasets using evaluation metrics.
Common metrics for language model bias include probability based metrics, which evaluate directly on token probabilities, and generation-based metrics, which evaluate on text outputs. Probability metrics include pseudo-log-likelihood (PLL), introduced for masked LMs by Nangia et al. (2020) and extended to autoregressive LMs by Felkner et al. (2023), (idealized) context association test (CAT/iCAT) Nadeem et al. (2021), all unmasked likelihood (AUL) Kaneko and Bollegala (2022), and the perplexity-based Language Model Bias (LMB) metric introduced by Barikeri et al. (2021). Generation metrics can be based on distributional similarity Bordia and Bowman (2019); Bommasani et al. (2023), auxiliary classifiers Gehman et al. (2020); Huang et al. (2020); Sheng et al. (2019), or hand-built lexicons of harmful words Nozza et al. (2021); Dhamala et al. (2021).
@@
4.2 Downstream Bias Evaluation
In addition to upstream metrics, there are many downstream methods for evaluation of social biases within a specific task or domain. Dedicated bias evaluation datasets exist for many NLP tasks, including machine translation Levy et al. (2021), coreference resolution Rudinger et al. (2018); Zhao et al. (2018), question answering Parrish et al. (2022), and sentiment classification Kiritchenko and Mohammad (2018); Mei et al. (2023). There are also domain-specific bias evaluations, such as Sap et al. (2019) in hate speech detection and Zhang et al. (2024) for the medical domain.
@@ 5 Discussion and Conclusion
Through detailed analysis of the behavior of NLI and TP bias evaluations at multiple levels (stereotype categories, specific stereotypes, and individual test instances) and across three model families, nine models, and three debiasing conditions for each, we find significant differences in bias evaluation results. First, we find that none of the metrics we tested to convert entailment probability tuples into percentile bias scores shows linear correlation with token probability bias scores. Second, we find that most of the NLI metrics we tested correlate poorly with each other, suggesting that NLI metrics are brittle in the context of coarse-grained aggregate percentile bias scores. Third, we show substantial differences in the behavior of raw per-instance bias scores for TP and NLI: TP is generally more sensitive to debiasing, while NLI bias evaluations are more likely to detect specific stereotype categories for which models are “underdebiased.” Finally, we show that both TP and NLI bias metrics are unexpectedly sensitive to the specific wording of counterstereotypical sentences, suggesting that the choice of counterfactual identities could be a source of noise in both types of bias evaluation. Because NLI and token probability show substantial differences in bias evaluation results, even on exactly the same set of bias definitions, we strongly recommend a combination of upstream, midstream, and downstream bias evaluations to ensure the most comprehensive bias audits of language models.
@@
There are also limitations specific to our WQ-NLI dataset. First, our dataset has extremely limited variation in template sentences, with almost all variety in the dataset coming from the predicates, identities, and names inserted in the templates. The second key limitation of WQ-NLI is the fact that the correct entailment prediction is always neutral. This paradigm follows prior work on NLI for bias evaluation (Dev et al., 2020). However, the correct labels in the MNLI training set are evenly split across entailment, contradiction, and neutral categories. Therefore, there is considerable difference in label distribution between the MNLI task finetuning dataset and the WQ-NLI evaluation dataset, which may have a negative impact on performance on the bias evaluation task.
Evaluation Limitations
[…]
@@
No generative AI or LLM system was used in ideation, experimentation, literature review, or writing. Coding assistants (Copilot and Gemini) were used to improve the styling of figures; however, the layout and content of figures was not AI-assisted. TeXGPT was also used to improve the typesetting of this paper.
@@ A.1 Predicate Categories
In order to facilitate a fine-grained analysis of TP and NLI as bias metrics, we manually sorted the attested harm predicates collected by Felkner et al. (2023) into eighteen categories. Categories are listed in Table 6. The attested harm predicate “are autistic” was included in the “mental illness” category, reflecting the context in which it is usually used as an anti-LGBTQ+ insult. However, the NIH considers autism a neurodevelopmental disorder, not a mental illness.
@@
A.2 Selected Category Results for Other Tested Models
Figs. 29 show the differences in log-probabilities for selected categories on TP (pink) and NLI (green) bias metrics. In general, NLI seems to be better at detecting “underdebiased” categories. Results are generally consistent across models for the naturalness/normalness category; however, the behavior is less clear in the drug use category for some models.
Figure 2: Entailment and token probability bias scores for selected categories on three BERT Base Uncased models (undebiased, news-debiased, and twitter-debiased). While results are less pronounced that BERT Large Uncased, NLI is still better at detecting underdebiased categories.
Figure 3: Entailment and token probability bias scores for selected categories on three BERT Base Cased models (undebiased, news-debiased, and twitter-debiased). Observed patterns are generally weaker but still present for cased BERT models.
Figure 4: Entailment and token probability bias scores for selected categories on three BERT Large Cased models (undebiased, news-debiased, and twitter-debiased). Observed patterns are generally weaker but still present for cased BERT models.
Figure 5: Entailment and token probability bias scores for selected categories on three RoBERTa Base models (undebiased, news-debiased, and twitter-debiased). For these models, NLI is clearly identifying lingering biases which TP scores do not.
Figure 6: Entailment and token probability bias scores for selected categories on three RoBERTa Large models (undebiased, news-debiased, and twitter-debiased). Observed pattern is very clear for naturalness/normalness category, but is not present for drug use category.
Figure 7: Entailment and token probability bias scores for selected categories on three GPT2 models (undebiased, news-debiased, and twitter-debiased). Observed pattern is very clear for naturalness/normalness category, but is not present for drug use category.
Figure 8: Entailment and token probability bias scores for selected categories on three GPT2 Medium models (undebiased, news-debiased, and twitter-debiased). Observed pattern is very clear for naturalness/normalness category, but is not present for drug use category.
Figure 9: Entailment and token probability bias scores for selected categories on three GPT2 XL models (undebiased, news-debiased, and twitter-debiased). Patterns are generally unclear, and the reason for the wide range of NLI bias scores on the naturalness/normalness category is unknown.

### v2

@@ 3.3 TP and NLI Results on Debiased Models
Table 5: Comparison of TP and NLI bias scores for raw (R.), news-debiased (N.), and Twitter-debiased (T.) models. Cases where NLI scores move in opposite direction to TP scores after debiasing are highlighted in bold. TP R., TP N., and TP T. columns are reproduced from results in Felkner et al. (2023).
@@ 3.4 Mutual Information Analysis
Until this point, we have been considering percentile bias scores that are calculated based on comparing the raw scores for the stereotypical and counterstereotypical sides of the WQ-TP/WQ-NLI datasets. In this section, we will consider the raw scores themselves. For token probability bias evaluation, bias scores are defined as in Nangia et al. (2020) and Felkner et al. (2023): the raw score for each test sentence is the sum of pseudo-log-probabilities for each of the tokens that are shared between the stereotypical and counterstereotypical test sentences. For NLI bias evaluation, the raw model outputs are tuples of six probabilities, as discussed above. For comparability, we consider log-probability versions of $M_{6}$ and $M_{9}$ from Table 2. We then consider the difference between stereotypical and counterstereotypical log-prob scores for each evaluation setup.
@@
Exploratory data analysis showed that log-probability differences are often very spread out, with large inter-quartile ranges and many outliers. To explain this high variability, we conduct manual analysis of the most extreme differences for a subset of models. We notice that a very small number of specific predicates dominate the most extreme examples; however, these predicates vary by model without a clear pattern across models. We also notice that several models seem to be very sensitive to the choice of counterfactual identity, with “cis” particularly overrepresented in highest-difference cases. This is likely because the tested models have seen relatively little training data using the word “cis,” and more data containing other counterfactual identities “straight,” “heterosexual,” and “cisgender.” In the analysis of extreme examples, we observe that the changes between raw and debiased models are largely as we expected. Debiasing reduces the magnitude of extreme differences and makes the extremes less dominated by specific predicates and counterfactuals.
@@
From these results, it is clear that there is considerable noise in bias scores (both percentile bias scores and raw log-probabilities) at both the per-instance and per-model levels. We thus conduct a more systematic analysis of which specific word choices have the largest effects on log-probability differences. We treat this as a feature selection problem, in which the potential features are binary columns corresponding to each possible model, finetuning condition, predicate, predicate category, name, pronoun, bias target group, counterfactual identity, and sentence template. We use mutual information regression, where these binary features predict a difference in log-probabilities, in order to determine the most sensitive, and thus meaningful, features.
@@
In this analysis, we expect to see higher mutual information for factors that should affect bias scores: model, finetuning condition, bias target groups, and predicate categories. We also expect that specific predicates will have some impact, but extremely high MI for certain predicates indicates that they may be introducing noise into the evaluation. For counterfactuals, names, pronouns, and templates, we expect to see very low MI values. Higher MI values for these categories indicate that incidental wording choices are introducing significant noise into the bias evaluation. The top ten MI factors for both TP and NLI evaluation are listed in Table 6.
Rk. TP Factor TP MI $M_{6}$ Factor $M_{6}$ MI $M_{9}$ Factor $M_{9}$ MI
1 Straight 0.029 GPT2 XL 0.096 GPT2 XL 0.077
2 raw 0.024 C: Nat./Norm. 0.016 C: Nat./Norm. 0.027
3 Heterosexual 0.024 Heterosexual 0.014 LGBTQ 0.024
4 twitter 0.021 LGBTQ 0.014 “are deviant” 0.017
5 Cisgender 0.011 “are deviant” 0.012 C: Lack of Belonging 0.017
6 GPT2 Med. 0.010 Queer 0.011 “are sexually deviant” 0.014
7 Cis 0.008 RoBERTa Base 0.010 Queer 0.013
8 No Predicate 0.008 Straight 0.010 “are not gay enough” 0.012
9 No Category 0.007 RoBERTa Large 0.009 C: Danger to Others 0.010
10 Is/And Template 0.005 GPT2 0.009 RoBERTa Base 0.009
@@
Table 6: Factors with highest mutual information values for token probability (left) and NLI (right). Specific wording of counterfactuals has a larger impact than expected; these instances are highlighted in bold. Stereotypes in quotes refer to specific attested harm predicates. "C:" refers to predicate categories, and "Nat./Norm." refers to the "naturalness/normalness" category.
@@
Our mutual information analysis shows that token probability bias evaluation seems to be somewhat more sensitive to choice of counterfactuals than NLI evaluation. $M_{6}$ is still sensitive to counterfactuals; $M_{9}$ seems to be less impacted by specific wording of counterfactuals. This result means that the wording of counterstereotypical sentences could unintentionally skew the resulting TP bias scores. NLI seems to be more sensitive to specific predicates and predicate categories. This means that NLI is likely better at detecting stereotypes where the model’s latent associations are strongest. However, the increased sensitivity to specific predicates indicates that wording choices of bias definitions are more likely to introduce noise into NLI bias evaluation. Overall, these results provide limited evidence that NLI may be more robust to counterfactual wording than TP, but they do not provide clear evidence that NLI is a more robust metric or generally “better” metric than TP.
@@ 4.1 Bias Measurement in LLMs
In this work, we use the definition of “bias” from Gallegos et al. (2024): “disparate treatment or outcomes between social groups that arise from historical and structural power asymmetries,” which includes both representational and allocational harms. Common metrics for language model bias include probability based metrics, which evaluate directly on token probabilities, and generation-based metrics, which evaluate on text outputs. Probability metrics include pseudo-log-likelihood (PLL), introduced for masked LMs by Nangia et al. (2020) and extended to autoregressive LMs by Felkner et al. (2023), (idealized) context association test (CAT/iCAT) Nadeem et al. (2021). Generation metrics can be based on distributional similarity Bordia and Bowman (2019); Bommasani et al. (2023), auxiliary classifiers Gehman et al. (2020); Huang et al. (2020); Sheng et al. (2019), or hand-built lexicons of harmful words Nozza et al. (2021); Dhamala et al. (2021).
@@ 5 Discussion and Conclusion
Through detailed analysis of the behavior of NLI and TP bias evaluations across seven model families, seventeen models, and three debiasing conditions, we find significant differences in bias evaluation results. First, we find that none of the metrics we tested to convert NLI probability tuples into percentile bias scores shows strong or moderate linear correlation with token probability bias scores. Second, we find that most of the NLI metrics we tested correlate poorly with each other, suggesting that NLI metrics are brittle in the context of coarse-grained aggregate percentile bias scores. Finally, we show that both TP and NLI bias metrics are unexpectedly sensitive to the specific wording of counterstereotypical sentences, suggesting that the choice of counterfactual identities could be a source of noise in both types of bias evaluation. NLI and token probability show substantial differences in bias evaluation results, even on exactly the same set of bias definitions, but there is no clear evidence than NLI is a better bias metric than TP. Therefore, we conclude that NLI is not viable as a replacement for token probability bias evaluation of langauge models.
@@
There are also limitations specific to our WQ-NLI dataset and our experiments. First, our dataset has extremely limited variation in template sentences, with almost all variety in the dataset coming from the predicates, identities, and names inserted in the templates. The second key limitation of WQ-NLI is the fact that the correct entailment prediction is always neutral. This paradigm follows prior work on NLI for bias evaluation (Dev et al., 2020). However, the correct labels in the MNLI training set are evenly split across entailment, contradiction, and neutral categories. Therefore, there is considerable difference in label distribution between the MNLI task finetuning dataset and the WQ-NLI evaluation dataset, which may have a negative impact on performance on the bias evaluation task. Finally, our evaluation is currently limited to open-weight models, though it may be extensible to closed-weight models with some modification.
@@
No generative AI or LLM system was used in ideation, experiment design, literature review, or writing. Coding assistants (Copilot and Gemini) were used to debug experiment and data analysis code and improve the styling of figures; however, the layout and content of figures was not AI-assisted. TeXGPT was also used to improve the typesetting of this paper.
@@ A.1 Predicate Categories
In order to facilitate a fine-grained analysis of TP and NLI as bias metrics, one author with relevant lived experience sorted the attested harm predicates collected by Felkner et al. (2023) into eighteen categories. Categories are listed in Table 7. The attested harm predicate “are autistic” was included in the “mental illness” category, reflecting the context in which it is usually used as an anti-LGBTQ+ insult. However, the NIH considers autism a neurodevelopmental disorder, not a mental illness.
@@
A.2 Detailed Model List
Table 8 lists all models studied in this work with citations and number of parameters. We used Llama 3.1 8B because the Llama 3.2 release does not include a model in the 7-8B size range.
Model Citation Params.
BERT Base Uncased Devlin et al. (2019) 110M
BERT Base Cased Devlin et al. (2019) 109M
BERT Large Uncased Devlin et al. (2019) 336M
BERT Large Cased Devlin et al. (2019) 335M
RoBERTa Base Liu et al. (2019) 125M
RoBERTa Large Liu et al. (2019) 561M
GPT2 Radford et al. (2019) 137M
GPT2 Medium Radford et al. (2019) 380M
GPT2 XL Radford et al. (2019) 1.61B
Llama 3.2 1B Dubey et al. (2024) 1.23B
Llama 3.2 3B Dubey et al. (2024) 3.21B
Llama 3.1 8B Dubey et al. (2024) 8B
Qwen 3 1.7B Team (2025) 1.7B
Qwen 3 4B Team (2025) 4.0B
Qwen 3 8B Team (2025) 8.2B
Gemma 7B Team et al. (2024) 7B
Mistral v0.3 7B Jiang et al. (2023) 7B
Table 8: Detailed listing of language models studied in our experiments. All are open-weight and available on HuggingFace.
A.3 Heatmap of Pairwise $R^{2}$ Values Among NLI Metrics
Fig. 1 shows the pairwise $R^{2}$ values for all combinations of tested NLI metrics. Most moderate and strong correlations correspond to metric pairs that are very similarly defined.
Figure 1: Pairwise $R^{2}$ values for NLI metrics. Strong and moderate correlations generally correspond to similarly formatted metrics.